# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

### 09 Jan 2024 v1

@@ Abstract
Human health can be critically affected by cardiovascular diseases, such as hypertension, arrhythmias, and stroke. Heart rate and blood pressure are important biometric information for the monitoring of cardiovascular system and early diagnosis of cardiovascular diseases. Existing methods for estimating the heart rate are based on electrocardiography and photoplethyomography, which require contacting the sensor to the skin surface. Moreover, catheter and cuff-based methods for measuring blood pressure cause inconvenience and have limited applicability. Therefore, in this thesis, we propose a vision-based method for estimating the heart rate and blood pressure. This thesis proposes a 2-stage deep learning framework consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers remote photoplethysmography (rPPG) signals for the acral and facial regions, and these phase-shifted rPPG signals are utilized to estimate the heart rate. In the second stage, BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP values. To improve the accuracy of estimating the heart rate, we employed a data augmentation method based on a frame interpolation model. Moreover, we designed BBP-Net to infer blood pressure within a predefined range by incorporating a scaled sigmoid function. To demonstrate the effectiveness of the proposed method, experiments were conducted on the MMSE-HR and V4V datasets. Our method resulted in estimating the heart rate with the mean absolute error (MAE) of 1.78 BPM, reducing the MAE by 34.31 % compared to the recent method, on the MMSE-HR dataset. The MAE for estimating the systolic blood pressure (SBP) and diastolic blood pressure (DBP) were 10.19 mmHg and 7.09 mmHg. On the V4V dataset, the MAE for the heart rate, SBP, and DBP were 3.83 BPM, 13.64 mmHg, and 9.4 mmHg, respectively. We expect that the proposed method will contribute to the advancement of sensing technology for the monitoring of human health.
@@ 1 Introduction
The cardiovascular system consists of the heart and blood vessels, and it circulates blood throughout the body, delivering oxygen and nutrients to tissues while removing waste substances. Abnormalities in the cardiovascular system can lead to various cardiovascular diseases such as cardiomyopathy, hypertensive heart disease, arrhythmias, with potentially severe implications for overall health. Moreover, cardiovascular diseases associated with blood vessels, such as hypertension and stroke, are contribute to increasing the global mortality rate (Gaidai et al., 2023). However, these diseases can be detected at an early stage through the monitoring of the heart rate and blood pressure using healthcare devices such as smartwatches. For example, arrhythmias caused by abnormalities in the sinoatrial node can be prevented through the measurement of bradycardia, tachyarrhythmia, and irregular heart rates. Additionally, consistent readings of systolic blood pressure (SBP) above 120 mmHg and diastolic blood pressure (DBP) above 90 mmHg may indicate the potential presence of hypertension. Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services (Etemadi et al., 2023; Aggarwal & Pandey, 2023) and estimation of biometric information (Alnaggar et al., 2023; Gupta et al., 2023).
@@
Heart rate and blood pressure are main biometric information for the monitoring of cardiovascular diseases (Perret-Guillaume et al., 2009; Fuchs & Whelton, 2020), and these information can be measured by utilizing various physiological sensors and medical equipment. Heart rate is commonly measured by using electrocardiography (ECG) and photoplethysmography (PPG). ECG uses electrodes attached to the body to record the electrical activity caused by contractions and relaxations of the heart. PPG is a non-invasive method that uses a light source and photodetector at the skin surface to measure the volumetric variations of blood in micro vessels. The heart rate is measured by computing peak-to-peak intervals in the time domain or by analyzing the power spectral density in the frequency domain from the ECG or PPG signals. On the other hand, blood pressure can be measured based on an oscillometric method, which records the magnitude of pressure oscillations using a cuff, or an invasive catheter-based approach. The catheter-based method involves the direct insertion of a sensor into an artery to measure continuous arterial blood pressure (ABP). Peak and valley values of the ABP signal are utilized as SBP and DBP, respectively. The blood pressure can also be estimated from pulse transit time (PTT) measured by ECG or PPG sensors at the heart and acral sites. PTT is the time taking for a blood pulse wave to travel from the heart to an acral site, and it known that PTT is closely correlated with blood pressure (Geddes et al., 1981). Although there have been proposed BP estimation methods by analyzing PTT from ECG or PPG signals (Barvik et al., 2021; Ganti et al., 2020), these approaches have the intrinsic limitation of requiring skin contact.
@@
Recently, camera sensors have been utilized to obtain physiological information, and it is called remote photoplethysmography (rPPG). From a facial video, the rPPG technique extracts subtle variations in skin color induced by cardiovascular pulse. The extracted temporal changes in pixel intensities are then transformed into continuous waveforms analogous to conventional PPG signals. The precision of the rPPG technique can be critically influenced by many factors such as ambient light conditions, motion artifacts, and different skin tones. Despite these challenges, rPPG receives much attention due to the non-contact and non-invasive attributes that facilitate remote monitoring of biometric information.
@@
Deep learning methods have been utilized to extract rPPG signals and to estimate the heart rate (Chen & McDuff, 2018; Yu et al., 2022; Ouzar et al., 2023). Additionally, various datasets (Bobbia et al., 2019; Niu et al., 2019; Soleymani et al., 2011; Li et al., 2018; Stricker et al., 2014; Zhang et al., 2016b; Revanur et al., 2021) have been released for the development and evaluation of these deep learning algorithms. PPG signals in the existing datasets were collected at acral sites such as fingertips, and therefore, there exist temporal discrepancies between the pulse waves in the PPG signals and facial videos. However, most previous methods ignore the time shift between the two modalities and adopt a time domain loss between PPG signals measured at an acral site and rPPG signals extracted from facial videos. For example, Lu et al. (2021) and Yu et al. (2022) employed negative Pearson correlation loss to measure the similarity between the PPG signals and estimated rPPG signals in the time domain. In this paper, we introduced two types of phase-shifted rPPG signals, referred to as acral rPPG and facial rPPG, to consider the temporal discrepancy between facial videos and PPG signals measured at an acral site. The acral rPPG is guided by a time domain loss function, while the facial rPPG is guided by frequency domain features which have temporally global characteristics.
@@
On the other hand, blood pressure can be estimated from PPG signals, both ECG and PPG signals, or facial videos. While PPG-based methods are widely utilized in wearable devices, this approach has the limitation of requiring physical contact to obtain PPG signals. Although the approach of utilizing both ECG and PPG signals can measure blood pressure based on PTT information, it also has the disadvantage of requiring contact with ECG and PPG sensors. Recently, camera-based methods have been proposed to estimate blood pressure in a non-contact manner by utilizing spatiotemporal features in facial videos. However, there is a performance gap between the camera-based and PPG-based methods. The objective of this study is to improve the performance of the camera-based approach by utilizing phase-shifted information in the acral and facial rPPG signals.
@@
This paper proposes a 2-stage deep learning pipeline consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers the acral and facial rPPG signals, and it is used to estimate the heart rate. In the second stage, BBP-Net integrates temporal features based on a multi-scale fusion (MSF) module and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP values. To improve the precision of blood pressure estimation, a scaled sigmoid function is employed in the BBP-Net to constrain the estimate into a certain range. Experiments were conducted on the MMSE-HR database (Zhang et al., 2016b) and V4V database (Revanur et al., 2021) to demonstrate the effectiveness of the proposed method for estimating heart rate and blood pressure from facial videos. The main contributions of this paper can be summarized as follows.
@@
    We employ a frame interpolation algorithm for the augmentation of bradycardia and pyknocardia video clips.
@@
    We introduce the concept of phase-shifted rPPG for considering the temporal discrepancy between pulse waves in facial videos and PPG signals.
@@ 2.1 Estimation of rPPG signals from facial videos
Image processing techniques and mathematical models have been utilized to analyze subtle variations in color intensities and to estimate rPPG signals. Poh et al. (2010) computed pulse waves by averaging on RGB color channels, and independent component analysis was conducted to extract rPPG signals. Lewandowska et al. (2011) estimated rPPG signals by conducting a channel selection process and principal component analysis on color intensities in a forehead region. De Haan & Jeanne (2013) proposed a chrominance-based method, called CHROM, to improve the performance of rPPG against motion artifacts. Wang et al. (2016) proposed the POS algorithm, which analyzes the plane orthogonal to the skin tone in the normalized RGB space. Meanwhile, with the success of deep learning methods for computer vision tasks, data driven approaches started to be applied in the estimation of rPPG signals.
The deep learning-based rPPG estimation technique focuses on learning trends in PPG that are robust to motion and external brightness changes from video sequence data with spatiotemporal characteristics. Chen & McDuff (2018) introduced a convolutional attention network (CAN) that employs a Siamese-structured 2 dimensional CNN model. This model takes both a difference map of adjacent viewpoints and the original viewpoint image. Spatial attentions are utilized to emphasize skin regions in the difference images, effectively integrating the models. Nowara et al. (2021) aimed to estimate spatial noise factors including motion and illumination using the inverse of spatial attentions of the CAN model. They integrated the LSTM to enhance spatial denoising and improve temporal robustness in PPG. Yu et al. (2019) proposed PhysNet, which employs a 3D CNN-based encoder to extract spatiotemporal features and a decoder along the time axis to estimate rPPG. They also introduced a loss function based on the Pearson correlation coefficient. Yu et al. (2022) proposed a loss function based on temporal difference convolution (TDC) using neighboring frame differences of a 4 dimensional feature map as query and key, and KL divergence of periodic component distributions. Yu et al. (2023) is a SlowFast Network for temporal improvement of the estimated signal, which has shown promising heart rate estimation accuracy even in cross-domain. However, existing studies primarily focuses on providing performance results of heart rate estimation using the power spectrum of rPPG, and less consideration is given to the blood pressure estimation steps based on rPPG.
@@ 2.2 Deep learning methods for estimating blood pressure
For non-invasive blood pressure estimation, researchers have attempted to develop deep learning algorithms utilizing physiological signal measured from various sensors. Miao et al. (2020) proposed a deep learning model based on ResNet and LSTM for continuous blood pressure estimation from single channel ECG. Panwar et al. (2020) introduced PP-Net, which consists of a 1D convolution and LSTM, to estimate SBP, DBP, and heart rate from PPG signals. Huang et al. (2022) utilized PTT information for blood pressure estimation and proposed a deep learning framework based on MLP-Mixer. Additionally, they introduced the multi-filter to multi-channel (MFMC) preprocessing of ECG and PPG signals. Similarly, Jia et al. (2022) estimate the blood pressure using a deep learning algorithm that learns PTT information from images of the face and palm. However, algorithms using sensor data as input are limited by their dependence on contact sensors, or the need to acquire data from multiple body parts, which restricts movement of subjects when measuring blood pressure.
Image-based blood pressure estimation methods have been conducted by combining rPPG estimation through mathematical modeling and leveraging deep learning model to estimate blood pressure from PPG signal. Schrumpf et al. (2021) conducted the study to estimate blood pressure from rPPGs by employing a rPPG extraction technique via a POS algorithm from ROI sequences and a PPG-based blood pressure estimation algorithm. Wu et al. (2022) also applied a POS algorithm to ROI sequence in 3 regions for rPPG extraction, and proposed a deep learning method to process rPPG and biometric information that can be estimated from it in parallel for blood pressure estimation. Bousefsaf et al. (2022) proposed an algorithm that extracts rPPG as G-channel averages of ROI, preprocesses them through continuous wavelet transform (CWT) in both time and frequency domains, and employs U-Net architecture to estimate continuous blood pressure signal. Chen et al. (2023) estimated blood pressure without rPPG extraction by employing spatiotemporal mapping of ROI sequences and a combined CNN and LSTM structure. While existing blood pressure estimation studies utilize mathematical modeling methods for rPPG extraction, we propose a deep learning algorithm for image-based rPPG estimation, which demonstrates improved performance in both heart rate and blood pressure estimation.
@@ 3.1 Overall training pipeline
Figure 1: Overview of the training pipeline of the proposed method.
This paper proposes a 2-stage deep learning framework which consists of DRP-Net and BBP-Net for estimating the heart rate and blood pressure from facial videos. The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP values. Fig. 1 presents the overview of the training pipeline of the proposed method. In the preprocessing step, facial regions are detected within a video clip to define a region of interest (ROI). The DRP-Net extracts acral and facial rPPG signals from the ROI sequence, and the BBP-Net predicts SBP and DBP values. To train the DRP-Net and BBP-Net, ground truth signals are generated from ABP signals. Pseudo PPG signals are computed by detrending and bandpass filtering to supervise the training of the DRP-Net Fast Fourier transform (FFT) and power spectrum density analysis are conducted to extract HR values from pseudo PPG and rPPG signals. Additionally, peaks and valleys are detected from the ABP signals to compute the ground truth SBP and DBP values. In Fig. 1, blue and red colors indicate ground truth and predicted signals, respectively.

### 06 Mar 2024 v2

@@ Abstract
Human health can be critically affected by cardiovascular diseases, such as hypertension, arrhythmias, and stroke. Heart rate (HR) and blood pressure (BP) are important physiological information for monitoring of cardiovascular system and early diagnosis of cardiovascular diseases. Previous methods for estimating heart rate are based on electrocardiography and photoplethysmography, which require contacting sensors to skin surface. Existing catheter and cuff-based methods for measuring blood pressure cause inconvenience and are difficult to be utilized in daily life. To address these limitations, this paper proposes a 2-stage deep learning framework for estimating heart rate and blood pressure from facial video. The proposed algorithm consists of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). DRP-Net infers remote photoplethysmography (rPPG) signals at acral and facial regions, and these phase-shifted rPPG signals are utilized to estimate heart rate. BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate systolic blood pressure (SBP) and diastolic blood pressure (DBP). We augmented facial video datasets in temporal aspect by utilizing a frame interpolation model to improve the accuracy of estimating heart rate. Moreover, we reduced the error in estimating blood pressure by incorporating a scaled sigmoid function in the BBP-Net. Experiments were conducted on MMSE-HR and V4V datasets to demonstrate the effectiveness of the proposed method. Our method achieved the state-of-the-art performance for estimating heart rate and blood pressure with significant margins compared to previous methods. Code is available at https://github.com/GyutaeHwang/phase_shifted_rPPG.
@@ 1 Introduction
The cardiovascular system consists of the heart and blood vessels, and it circulates blood throughout the body, delivering oxygen and nutrients to tissues while removing waste substances. Abnormalities in the cardiovascular system can lead to various cardiovascular diseases such as cardiomyopathy, hypertensive heart disease, arrhythmias, with potentially severe implications for overall health. Moreover, cardiovascular diseases associated with blood vessels, such as hypertension and stroke, are contribute to increasing the global mortality rate (Gaidai et al., 2023). However, these diseases can be detected at an early stage through the monitoring of the heart rate (HR) and blood pressure (BP) using healthcare devices such as smartwatches. For example, arrhythmias caused by abnormalities in the sinoatrial node can be prevented through the measurement of bradycardia, tachyarrhythmia, and irregular heart rates. Additionally, consistent readings of systolic blood pressure (SBP) above 120 mmHg and diastolic blood pressure (DBP) above 90 mmHg may indicate the potential presence of hypertension. Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services (Etemadi et al., 2023; Aggarwal & Pandey, 2023) and estimation of physiological information (Alnaggar et al., 2023; Gupta et al., 2023).
@@
Heart rate and blood pressure are main physiological information for the monitoring of cardiovascular diseases (Perret-Guillaume et al., 2009; Fuchs & Whelton, 2020), and these information can be measured by utilizing various physiological sensors and medical equipment. Heart rate is commonly measured by utilizing electrocardiography (ECG) and photoplethysmography (PPG). ECG employs electrodes attached to the body to record the electrical activity caused by contractions and relaxations of the heart. PPG is a non-invasive method that uses a light source and photodetector attached at skin surfaces to measure the volumetric variations of blood in micro vessels. The heart rate can be measured by computing peak-to-peak intervals in the time domain or by analyzing power spectrum in the frequency domain from ECG or PPG signals. On the other hand, blood pressure can be measured based on an oscillometric method, which records the magnitude of oscillations using a blood pressure cuff. A catheter-based method is an invasive approach to measure blood pressure, and it involves direct insertion of a sensor into an artery to measure real-time arterial blood pressure (ABP). SBP and DBP can be computed from peak and valley values of the ABP signals. Pulse transit time (PTT) is temporal delay of a blood pulse wave to travel from the heart to an acral site such as a fingertip, and it is known that PTT is closely correlated with blood pressure (Geddes et al., 1981). Although there have been proposed blood pressure estimation methods by analyzing PTT from ECG or PPG signals (Barvik et al., 2021; Ganti et al., 2020), these approaches have the intrinsic limitation of requiring skin contacts.
@@
Recently, camera sensors have been utilized to obtain physiological signals, and it is called remote photoplethysmography (rPPG). From a facial video, the rPPG technique extracts subtle variations in skin color induced by cardiac pulse. The extracted temporal changes in pixel intensities are then transformed into continuous waveforms analogous to conventional PPG signals. The precision of the rPPG technique can be critically influenced by many factors such as ambient light conditions, motion artifacts, and different skin tones. Despite these challenges, rPPG receives much attention due to the non-contact and non-invasive attributes that facilitate remote monitoring of physiological information.
@@
Deep learning methods have been utilized to extract rPPG signals and to estimate the heart rate (Chen & McDuff, 2018; Yu et al., 2022; Ouzar et al., 2023). Additionally, various datasets (Bobbia et al., 2019; Niu et al., 2019; Soleymani et al., 2011; Li et al., 2018; Stricker et al., 2014; Zhang et al., 2016b; Revanur et al., 2021) have been released for the development and evaluation of these deep learning algorithms. PPG signals in the existing datasets were collected at acral sites such as fingertips, and therefore, there exist temporal discrepancies between the pulse waves in the PPG signals and facial videos. However, most previous methods neglect the time shift between the two modalities and adopt a time domain loss between PPG signals measured at an acral site and rPPG signals extracted from facial videos. For example, Lu et al. (2021) and Yu et al. (2022) employed negative Pearson correlation loss to measure the similarity between the PPG signals and estimated rPPG signals in the time domain. In this paper, we introduced two types of phase-shifted rPPG signals, referred to as acral rPPG and facial rPPG, to consider the temporal discrepancy between facial videos and PPG signals measured at an acral site. The acral rPPG signal is guided by a time domain loss function, while the facial rPPG signal is guided by frequency domain features which have temporally global characteristics.
@@
On the other hand, blood pressure can be estimated from PPG signals, both ECG and PPG signals, or facial videos. While PPG-based methods are widely utilized in wearable devices, this approach has the limitation of requiring physical contact to obtain PPG signals. Although the approach of utilizing both ECG and PPG signals can measure blood pressure based on PTT, it also has the disadvantage of requiring contact with ECG and PPG sensors. Recently, camera-based methods have been proposed to estimate blood pressure in a non-contact manner by utilizing spatiotemporal features in facial videos. However, there is a performance gap between the camera-based and PPG-based methods. The objective of this study is to improve the performance of the camera-based approach by utilizing phase-shifted information in the acral and facial rPPG signals.
@@
This paper proposes a 2-stage deep learning pipeline consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers the acral and facial rPPG signals, and it is used to estimate the heart rate. In the second stage, BBP-Net integrates temporal features based on a multi-scale fusion (MSF) module and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP. To improve the precision of blood pressure estimation, a scaled sigmoid function is employed in the BBP-Net to constrain the estimate into a certain range. Experiments were conducted on the MMSE-HR database (Zhang et al., 2016b) and V4V database (Revanur et al., 2021) to demonstrate the effectiveness of the proposed method for estimating heart rate and blood pressure from facial videos. The main contributions of this paper can be summarized as follows.
@@
    We employ a frame interpolation algorithm for the augmentation of bradycardia and tachycardia video clips.
@@
    We introduce the concept of phase-shifted rPPG signals for considering the temporal discrepancy between pulse waves in facial videos and PPG signals.
@@ 2.1 Estimation of rPPG signals from facial videos
The conventional rPPG extraction method analyzed subtle variations in color intensities using image processing techniques and mathematical modeling. Poh et al. (2010) computed pulse waves by averaging RGB color channels, and independent component analysis was conducted to extract rPPG signals. On the other hand, Lewandowska et al. (2011) estimated rPPG signals by conducting a channel selection process and principal component analysis on color intensities in the forehead region. De Haan & Jeanne (2013) proposed a chrominance-based method, called CHROM, to improve the performance of extracting rPPG signals against motion artifacts. Wang et al. (2016) proposed the POS algorithm, which analyzes the plane orthogonal to the skin tone in the normalized RGB space. Meanwhile, the success of deep learning methods for computer vision tasks, data driven approaches began to be applied in the estimation of rPPG signals.
The deep learning-based rPPG estimation technique focuses on learning trends of PPG signals that are strongly reactive to motion and external brightness changes from video sequence data with spatiotemporal characteristics. Chen & McDuff (2018) introduced the convolutional attention network (CAN) that employs a Siamese-structured two-dimensional convolutional neural network (CNN) model. This model takes both the original temporal point image and a difference map of adjacent points. Spatial attentions are utilized to emphasize skin regions in the difference images, effectively integrating the models. Nowara et al. (2021) aimed to estimate spatial noise factors including motion and illumination using the inverse of spatial attentions of the CAN model. They integrated the LSTM to enhance spatial denoising and improve temporal robustness in PPG signals. Yu et al. (2019) proposed PhysNet3D, which employs a three-dimensional CNN-based encoder to extract spatiotemporal features and a deconvolution of the time axis to estimate rPPG signals. They also introduced a loss function based on the Pearson correlation coefficient. Yu et al. (2022) proposed a video transformer based on temporal difference convolution (TDC), which utilizes adjacent frame differences of a 4-dimensional feature map as query and key. They also proposed a KL divergence-based loss function for periodic component distributions. Yu et al. (2023) utilized a SlowFast Network to improve the temporal representation of the estimated signal, which has demonstrated promising accuracy in heart rate estimation, even in cross-domain experiments. However, existing studies primarily focus on providing performance results of heart rate estimation using the power spectrum of rPPG signals. Therefore, scholars have given less consideration to the blood pressure estimation steps based on rPPG signals.
@@ 2.2 Deep learning methods for estimating blood pressure
Non-invasive blood pressure estimation methods aim to develop deep learning algorithms utilizing physiological signals measured from user-friendly sensors. Miao et al. (2020) proposed a deep learning model based on ResNet and Long Short-Term Memory (LSTM) for continuous blood pressure estimation from single channel ECG signal. Panwar et al. (2020) introduced PP-Net, which consists of a one-dimensional convolution blocks and LSTM, to estimate SBP, DBP, and heart rate from PPG signals. Huang et al. (2022) utilized the concept of PTT for blood pressure estimation and proposed a deep learning framework based on MLP-Mixer. Additionally, the researchers introduced the multi-filter to multi-channel (MFMC) technique for preprocessing ECG and PPG signals. Similarly, Jia et al. (2022) estimate blood pressure using a proposed deep learning model that analyzes PTT from image sequences of the face and palm. However, the mentioned models require data to be acquired from contact sensors or multiple physiological sites, which have limitations that restrict the user’s movements.
Image-based blood pressure estimation methods have been developed by combining conventional rPPG extraction methods with deep learning models that estimate blood pressure from PPG signals. Schrumpf et al. (2021) conducted a study to estimate blood pressure from rPPG signals by employing a rPPG extraction technique via a POS algorithm from ROI sequences and a PPG-based blood pressure estimation model. Wu et al. (2022) also applied a POS algorithm to three ROI sequences for rPPG extraction and proposed FS-Net, which processes rPPG signals and physiological information in parallel for blood pressure estimation. Bousefsaf et al. (2022) used green channel average for rPPG extraction and proposed a method to restore continuous ABP signal based on continuous wavelet transform and U-Net. Chen et al. (2023) estimated blood pressure without rPPG extraction by employing spatiotemporal mapping of ROI sequences and a combined CNN and LSTM structure. While existing blood pressure estimation studies utilize non-parametric methods for rPPG extraction, we propose a deep learning algorithm for image-based rPPG estimation, which demonstrates improved performance in both heart rate and blood pressure estimation.
@@ 3.1 Overall training pipeline
This paper proposes a two-stage deep learning framework consisting of DRP-Net and BBP-Net to estimate heart rate and blood pressure from facial videos. The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP. Fig. 1 presents the overview of the training pipeline of the proposed method. In the preprocessing step, facial regions are detected within a video clip to define a region of interest (ROI). The DRP-Net learns spatiotemporal features from the ROI sequence and extracts acral and facial rPPG signals via dual prediction heads. Afterwards, The BBP-Net learns temporal multi-scale information from phase-shifted rPPG signals to predict SBP and DBP. To generate ground truth data, ABP signals are utilized to calculate pseudo PPG signals, heart rate, SBP and DBP.
Figure 1: Overview of the training pipeline of the proposed method. Blue and red colors indicate ground truth and predicted physiological information, respectively.